OpenStudy (anonymous):

A sandpit in the shape of a rectangular prism has length 7 feet, width 5 feet, and height 1.75 feet. It is filled to the brim with sand. Joe puts this sand into a second sandpit having the same shape but a larger base. He needs 17.5 cubic feet of sand to fill the extra space inside the second sandpit. If the height of the two sandpits is the same, what are the dimensions of the base of the second sandpit? Answer 5 feet x 2 feet 10 feet x 7 feet 9 feet x 5 feet 20 feet x 25 feet

OpenStudy (anonymous):

volume of first one is 7*5*1.75=61.25 cubic feet and it needs 17.5 cubic feet extra sand so, the volume of second one is 61.25+17.5=78.75 cubic feet is has same height as first one so the area of base is 78.75/1.75=45 which only satisfy 9 feet x 5 feet the third selection

OpenStudy (anonymous):

thank-you!!